OpenStudy (anonymous):

Sketch the Asymptotes and graph the function of y=5/(x-2)+3 Help an fellow math student! please!

OpenStudy (welshfella):

the asymptote occurs when the denominator of the fraction = 0

OpenStudy (anonymous):

so -2, -3? or is does the +3 just move the graph over?

OpenStudy (welshfella):

What value of x makes x - 2 = 0?

OpenStudy (anonymous):

-2

OpenStudy (anonymous):

i mean 2

OpenStudy (welshfella):

thats correct theres a vertical asymptote at x = 2

OpenStudy (anonymous):

So how would i determine the graph? enter in 5/(x-2) and move it over 3

OpenStudy (welshfella):

first think about the value of the function when x approaches infinity and negative infinity the fraction will approach zero and the function will approach 3 , right?

OpenStudy (anonymous):

Could you clarify for a layman? haha

OpenStudy (welshfella):

if x is say 2 million then 5 /(x + 2) is pretty near to zero. Right?

OpenStudy (welshfella):

- though it never reaches zero

OpenStudy (welshfella):

- similarly for x = - 2 million and because of the +3 that means that the function approaches 3 So there is a horizontal asymptote at y=3

OpenStudy (anonymous):

So would the horizontal asymptote be 3?

OpenStudy (anonymous):

haha i found it correctly then

OpenStudy (welshfella):

y=3 yes next thging to find is the y ntercept Put x = 0 and find value of y: y = 5 /(0-2) + 3 = -2.5 + 3 = 0.5 so the graph cuts the y axis at y = 0.5

OpenStudy (welshfella):

it will look something like thuis